OpenStudy (anonymous):

find dx/dt by implicit differentiation: x-sqrt(t) = tx PLEASE HELP ME! AND EXPLAIN Step bY Step sO I CAN UNDERSTAND! THANK YOU!

ganeshie8 (ganeshie8):

basically you need to find x'

ganeshie8 (ganeshie8):

t is the variable you differentiating with respect to

ganeshie8 (ganeshie8):

x-sqrt(t) = tx differentiate both sides

ganeshie8 (ganeshie8):

d/dt [x-sqrt(t)] = d/dt [tx]

ganeshie8 (ganeshie8):

x' - 1/2sqrt(t) = d/dt [tx]

ganeshie8 (ganeshie8):

right hand side you must use 'product rule'

ganeshie8 (ganeshie8):

can you take this from here ?

OpenStudy (anonymous):

after i take the product rule on the right side, then is that it?

ganeshie8 (ganeshie8):

wat do you get after taking product rule ?

ganeshie8 (ganeshie8):

you need to solve for x' in the end, and then thats it :)

OpenStudy (anonymous):

so it would be d/dt (t)'x + (t)(x)' right?

OpenStudy (anonymous):

what would be the derivative of t???

ganeshie8 (ganeshie8):

derivative of 't' with respect to 't' is 1 d -- (t) = 1 dt

OpenStudy (anonymous):

just like x?

OpenStudy (anonymous):

so my right side answer would be just x+t?

ganeshie8 (ganeshie8):

yes ! derivative of 'x' with respect to 'x' is 1

ganeshie8 (ganeshie8):

not exactly, derivative of 'x' with respect to 't' is NOT 1

ganeshie8 (ganeshie8):

let me show the complete solution

ganeshie8 (ganeshie8):

x-sqrt(t) = tx differentiate both sides with respect to t d/dt[x-sqrt(t)] = d/dt[tx] x' - 1/2sqrt(t) = d/dt [tx] x' - 1/2sqrt(t) = tx' + x solve for x' now

ganeshie8 (ganeshie8):

x' - 1/2sqrt(t) = tx' + x x'(1-t) = x + 1/2sqrt(t) x + 1/2sqrt(t) x' = ------------ 1-t

ganeshie8 (ganeshie8):

you may simplify if you want, else you can leave it as it as

ganeshie8 (ganeshie8):

btw, x' is the dx/dt

OpenStudy (anonymous):

Thank you so much!

ganeshie8 (ganeshie8):

np :)